OpenStudy (anonymous):

Integral of (2-v)/(v^2-4v+1) dv

OpenStudy (anonymous):

try \(u=v^2-4v+1, du=(2v-4)dv,-\frac{1}{2}du=(2-v)dv\) and you should be good to go

OpenStudy (anonymous):

Cool, I'll try it. Thanks for the direction. I'll let you know how it goes...

OpenStudy (anonymous):

Works. Thx..